OpenStudy (anonymous):

After one month at a new job you decide to save $50. Each month you plan to increase your savings by five dollars. Write an explicit formula to model the amount you will save each month. How much will you save in the seventh month?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

x = number of months, y = number of months after the first month MS = Money saved after x months The first month in the problem is merely x($50) = 1(50) = $50 the first month Thereafter, he increases each month by $5, so: MS = x($50) + y($5) So after two months he has: MS = 2($50) + 1($5)...remember, the first month he did not save an additional $5. We might want to re-write this as: MS = x($50) + (x - 1)($5) instead of using y, so after two months he has saved a total of: MS = 2($50) + (2 - 1)($5) = $200 + $5 = $205 In 6 months he will have saved: MS = 6($50) + (6 - 1)($5) = $300 + $5(5) = $300 + $25 = $325 The equation thus should be: MS = x($50) + (x - 1)($5) ---------------------------------------… Seventh month: MS = 7($50) + (7 - 1)($5) = $350 + $30 = $380 ---------------------------------------… This is the same kind of problem, only this time, in percents D = Diamond Value x = number of years D = $3500 + [(0.05)(x)*$3500)] D = $3500 + [(0.05)(6)*($3500)] D = $3500 + [0.3*$3500] D = $3500 + $1050 D = $4550 Remember that percents have to be in decimal percentage! so 5% = 5/100 = 0.05

OpenStudy (stamp):

s = 50 + 5m s = savings m = months